# On Partition Functions for Time-Inhomogeneous Branching Random Walks

### 15 Mar 2026 v1

@@ 4.3 Other Related Models
In this final section, we would other related models with time-inhomogeneous BRWs discussed in the present paper..
An interesting implication of Theorem 3.3 is that if we embed the (binary) BRW into the unit interval $[0,1]$, the covariance structure $K(x,y)$ (for $x,y\in[0,1]$) differs from that of GMC (which takes the form $\log|x-y|+O(1)$): it is not log-correlated. The universality behavior we established enables us to investigate whether we can construct a GMC-type object beyond log-correlated fields, which defines a variant of GMC measure. In this setting, the covariance kernel takes form of $F(x,y)+O(1)$, where $F$ is a positive, strictly non-increasing function satisfying $F(x,y)<\log|x-y|$ for all pairs of $x,y$.
Additionally, the results regarding to time-inhomogeneity of the increments discussed in the present paper raises a more fundamental questions: does addition of environmental fluctuations impact phase transitions (or other qualitative behavior) in disordered systems?
@@
On the other hand, the model of BRWs has recently been viewed as a specific example of directed polymers on infinite graphs (see [11] for an introduction to this topic), and we hope our results can be extended to other related models, such as $\lambda$-biased BRWs on Galton-Watson trees. Finally, it is also interesting to extend the results to environments that are not independent: i.e., time-correlated random fields. There are more interesting dependence of time, compared to monotonic and macroscopic decay in variance discussed in the present paper in the modern literature. For example, the study of the influence of environmental fluctuations, or stochastic impurities, on directed polymers has been initiated in [7]. The authors sincerely hope that more interesting properties and phenomena will be discovered in the time-inhomogeneous settings of these models.

### 27 Mar 2026 v2

@@ 4.4 Other Related Models
On the other hand, the model of BRWs has recently been viewed as a specific example of directed polymers on infinite graphs (see [13] for an introduction to this topic), and we hope our results can be extended to other related models, such as $\lambda$-biased BRWs on Galton-Watson trees. Finally, it is also interesting to extend the results to environments that are not independent: i.e., time-correlated random fields. There are more interesting dependence of time, compared to macroscopic decay in variance discussed in the present paper in the modern literature. For example, the study of the influence of environmental fluctuations, or stochastic impurities, on directed polymers has been initiated in [9]. The authors sincerely hope that more interesting properties and phenomena will be discovered in the time-inhomogeneous settings of these models.